Give the standard form of the equation of the parabola with the given characteristics. Vertex: (-3,1) Focus: (-1,1)
Answer by DSMLMD(16)

Vertex (-3,1) and Focus at (-1,1)

Because the y-vertex and y-focus is same, the equation of parabola is:
(y - b)^2 = 4p(x - a)

Vertex:
(-3,1)
(a,b)

Focus Point:
(-1,1)
((p+a),b)

If
b = 1
a = -3
p + a = -1

then,
p + a
-1 = p + -3
2 = p

the parabola equation is:
(y - b)^2 = 4p(x - a)
(y - 1)^2 = 4(2)(x - (-3))
(y - 1)^2 = 8(x + 3)